# definition

### v4

@@ Appendix D Proof of main result (Theorem 1)
Here recall that $\bm{x}$ is an $N\times\lvert\mathbb{S}\rvert$ matrix whose $i$th row, $x_{i}$, describes the state of arm $i$; $P_{i}$ is the transition probability matrix for arm $i$ under the optimal single-armed policy; $\gamma=\exp(-1/(2\tau))$ for $\tau$ defined in 1. To build intuition for $h(\bm{x},D)$, consider the term corresponding to $g=r^{*}$ and $\ell=0$. In this case, the term $\sum_{i\in D}\left\langle x_{i}-\mu_{i}^{*},g_{i}\right\rangle$ measures the difference between the reward obtained by applying the optimal single-armed policies to arms in $D$ and the reward upper bound given by the LP relaxation. A similar interpretation holds for the differences in costs.
In Lemma 3 below, we show that $h(\bm{x},D)$ is well-defined and establish its two key properties, which play a critical role in our analysis. The proof of Lemma 3 is given in Section E.2.
@@
We are now ready to define the function $h_{\textnormal{ID}}(\cdot,\cdot)$ used to construct the Lyapunov function $V$. For any system state $\bm{x}$ and $m\in[0,1]_{N}$ (where recall that $[0,1]_{N}$ is the set of integer multiples of $1/N$ within the interval $[0,1]$), $h_{\textnormal{ID}}(\bm{x},m)$ is defined as
@@
That is, $h_{\textnormal{ID}}(\bm{x},m)$ is an upper envelope of the subset Lyapunov functions $h(\bm{x},[Nm^{\prime}])$’s. The function $h_{\textnormal{ID}}(\bm{x},m)$ has properties similar to those in Lemma 3, which we state as Lemma 7 and prove in Appendix E.3.
@@
We next introduce the concept of the focus set, which is directly tied to the function $m(\cdot)$ in the Lyapunov function $V$. The focus set is a dynamic subset of arms based on the current system state. Specifically, for any system state $\bm{x}$, the focus set is defined as the set $[Nm(\bm{x})]$, where $m(\bm{x})$ is given by
@@
The focus set is introduced because it has several desirable properties that are useful for the analysis. First, under the ID policy, almost all the arms in the focus set, except for $O(\sqrt{N})$ arms, can follow the optimal single-armed policies. Additionally, as the focus set evolves over time, it is almost non-shrinking, and its size is closely related to the value of the function $h_{\textnormal{ID}}(\cdot,\cdot)$. These properties are formalized as Lemmas 8, 9 and 10, which are presented in Appendix F.
@@
where $L_{h}$ is the Lipschitz constant in Lemma 3 and $K_{\text{conf}}$ is the positive constant in Lemma 8.
@@
The proofs of Lemmas 4 and 5 are provided in Appendix G. It is then straightforward to combine these two lemmas to get Theorem 1.
Appendix E Lemmas and proofs for subset Lyapunov functions
In this section, we prove lemmas for the subset Lyapunov functions $h(\bm{x},D)$ and $h_{\textnormal{ID}}(\bm{x},m)$, defined in (22) and (26). In Section E.1, we prove a preliminary lemma that helps us utilize 1. In Section E.2, we prove Lemma 3, which addresses the properties of the subset Lyapunov functions $h(\bm{x},D)$. Finally, in Section E.3, we present and prove Lemma 7, which establishes properties of the function $h_{\textnormal{ID}}(\bm{x},m)$.
E.1 Lemma for utilizing mixing time bound of transition matrices
Recall that in 1, we assume that for each arm $i$, the mixing times of the transition matrix $P_{i}$ under the optimal single-armed policy ${{\bar{\pi}}^{*}}$ is uniformly bounded by a constant $\tau$. In this section, we prove a lemma that allows us to utilize this assumption.
Let $\Xi_{i}$ denote a matrix whose rows are the same vector $\mu_{i}^{*}$, where recall that $\mu_{i}^{*}$ is the stationary distribution of $P_{i}$. Lemma 6 stated below bounds the infinite series $\sum_{\ell=0}^{\infty}\max_{i\in[N]}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\gamma^{-\ell}$ by a constant independent of $N$.
@@ E.1 Lemma for utilizing mixing time bound of transition matrices
Note that Lemma 6 is the only place in our proofs where 1 is directly invoked; all of our later proofs utilize 1 indirectly by invoking Lemma 6.
Proof of Lemma 6.
@@
Moreover, by 1, it is not hard to see that
@@
Observe that $\tau\geq 1$, 0<1/($20<1/(2\tau)\leq 1/$ 2, and $\exp(-a)\leq 1-(1-1/\sqrt{e})a$ for any $20<a\leq 1/$ 2, so
@@
E.2 Proof of Lemma 3
In this subsection, we prove Lemma 3, which is about properties of the Lyapunov function $h(\bm{x},D)$:
@@ E.2 Proof of Lemma 3
Lemma 3 is restated as follows. See 3
In the proof of Lemma 3, we will frequently use the following equivalent form of $h({\bm{X}},D)$:
@@
By Lemma 6, $\sum_{\ell=0}^{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}/\gamma^{\ell}$ is finite, so the expression in (35) is also finite. Taking maximum over $g\in\mathcal{G}$, because $\mathcal{G}$ is a finite set, we have
$h(\bm{x},D)=\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\sum_{i\in D}\left\langle(x_{i}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\right\rvert<\infty.$
@@
Next, we bound the term $\mathbb{E}\left[\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$. Let $\epsilon_{i,t}\in\mathbb{R}^{\mathbb{S}}$ be the random vector given by $\epsilon_{i,t}=X_{i,t+1}-X_{i,t}P_{i}$ for $i\in D$. Then for each state $s\in\mathbb{S}$, $\epsilon_{i,t}(s)$ conditioned on ${\bm{X}}_{t}$ is a Bernoulli distribution with mean 00. Consequently,
$\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert$
@@
where the inequality in (44) is due to $\sum_{\ell=0}^{\infty}\max_{i\in D}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}/\gamma^{\ell}\leq C_{\tau}$ (Lemma 6).
@@
E.3 Properties of $h_{\textnormal{ID}}(\cdot,\cdot)$
@@
    where $L_{h}>0$ is the Lipschitz constant given in Lemma 3.
@@
    where $C_{h}>0$ is the constant given in Lemma 3.
@@
where (48) is due to the Lipschitz continuity of $h(x,D)$ with respect to $D$, as established in Lemma 3. We have thus proved (46).
@@
$\mathbb{E}\left[(h_{\textnormal{ID}}({\bm{X}}_{t+1},m)-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m))^{+}\;\middle|\;{\bm{X}}_{t}\right]$
@@
$\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left\lvert\sum_{i\in[Nm^{\prime}]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert^{2}\;\middle|\;{\bm{X}}_{t}\right]^{1/2}$
@@
Appendix F Lemmas for focus set
In this section, we present and prove three lemmas about properties of the focus set. Recall that for any system state $\bm{x}$, the focus set is defined as the set $[Nm(\bm{x})]$, where $m(\bm{x})$ is given by
@@ Appendix F Lemmas for focus set
$\frac{1}{N}\mathbb{E}[\left(Nm({\bm{X}}_{t})-N_{t}^{*}\right)^{+}\;|\;{\bm{X}}_{t}]\leq\frac{K_{\text{conf}}}{\sqrt{N}},\quad\text{with probability }1,$
@@
$1-m({\bm{X}}_{t})\leq\frac{1}{\eta_{c}N}h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))+\frac{K_{\text{cov}}}{N},\quad\text{with probability}1,$
@@
F.1 Proof of Lemma 8 (Majority conformity)
@@
where (F.1) follows from the strict slope of the remaining budget $\overline{C}^{*}([Nm])$ (Lemma 2). Therefore, choosing $m=m({\bm{X}}_{t})-(\Delta_{t}+M_{c})/(\eta_{c}N)$, we obtain
@@
F.2 Proof of Lemma 9 (Almost non-shrinking)
@@
where $C_{h}>0$ is the positive constant given in Lemma 3.
@@
We upper bound $\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]$ by coupling ${\bm{X}}_{t+1}$ with a random element ${\bm{X}}_{t+1}^{\prime}$ constructed as follows: Let ${\bm{X}}_{t+1}^{\prime}$ be the system state at step $t+1$ if we were able to set $A_{i,t}=\widehat{A}_{i,t}$ for all $i\in[N]$. From the drift condition of the Lyapunov function $h_{\textnormal{ID}}(\cdot,D)$, as established in Lemma 7, we obtain:
@@
$\mathbb{E}\left[\big{(}h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\big{)}^{+}-\big{(}h_{\textnormal{ID}}({\bm{X}}_{t+1}^{\prime},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\big{)}^{+}\;\middle|\;{\bm{X}}_{t}\right]$
@@
where we have applied Lemma 8 to bound the expression in (65).
@@
where the right-hand side can be further bounded using Lemma 11, which states that
@@
F.3 Proof of Lemma 10 (Sufficient coverage)
@@
Appendix G Proofs of Lemma 4 and Lemma 5
In this section, we provide two final lemmas, Lemma 4 and Lemma 5, which together imply Theorem 1. We prove these two lemmas in Sections G.1 and G.2, respectively.
G.1 Proof of Lemma 4
@@
where (72) is due to the definition of $N_{t}^{*}$; (73) is due to the bound on $\mathbb{E}[\left(Nm({\bm{X}}_{t})-N_{t}^{*}\right)^{+}]$ in Lemma 8, and (74) directly follows from the definition of $h_{\textnormal{ID}}(\bm{x},1)$. We further bound the first two expressions in (74) in terms of $V({\bm{X}}_{t})$ as follows:
@@ G.2 Proof of Lemma 5
where the last inequality follows from Lemma 9. To bound $\mathbb{E}\left[h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))\mid{\bm{X}}_{t}\right]$, observe that by Lemma 8, all but $O(\sqrt{N})$ arms in $[Nm({\bm{X}}_{t})]$ follow the optimal single-armed policies, so the drift condition of $h_{\textnormal{ID}}$ applies to this set of arms. As formalized in Lemma 11, we can thus show that
@@
$+\left(2C_{h}+C_{h}K_{\text{conf}}+L_{h}K_{\text{mono}}\right)\sqrt{N}.$
To further bound $h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))$ in (79) in terms of $V({\bm{X}}_{t})$, we apply Lemma 10 to get:
@@
Therefore, we can bound the long-run-averaged expectation of $\mathbb{E}\left[V(\bm{x})\right]$ as
@@
Finally, we give a more explicit form of the constant $C_{\textnormal{ID}}$ in Theorem 1:
@@
  • If $c_{\max}\geq\alpha_{\min}$, we combine the optimality gap bound in Lemma 4 with (82) to get

### v5

@@ Appendix E Proof of main result (Theorem 1)
Here recall that $\bm{x}$ is an $N\times\lvert\mathbb{S}\rvert$ matrix whose $i$th row, $x_{i}$, describes the state of arm $i$; $P_{i}$ is the transition probability matrix for arm $i$ under the optimal single-armed policy; $\gamma=\exp(-1/(2\tau))$ for $\tau$ defined in ˜1. To build intuition for $h(\bm{x},D)$, consider the term corresponding to $g=r^{*}$ and $\ell=0$. In this case, the term $\sum_{i\in D}\left\langle x_{i}-\mu_{i}^{*},g_{i}\right\rangle$ measures the difference between the reward obtained by applying the optimal single-armed policies to arms in $D$ and the reward upper bound given by the LP relaxation. A similar interpretation holds for the differences in costs.
In Lemma 3 below, we show that $h(\bm{x},D)$ is well-defined and establish its two key properties, which play a critical role in our analysis. The proof of Lemma˜3 is given in Section˜F.2.
@@
We are now ready to define the function $h_{\textnormal{ID}}(\cdot,\cdot)$ used to construct the Lyapunov function $V$. For any system state $\bm{x}$ and $m\in[0,1]_{N}$ (where recall that $0,1]_{N}$ is the set of integer multiples of $1/N$ within the interval $[0,1]$), $h_{\textnormal{ID}}(\bm{x},m)$ is defined as
@@
That is, $h_{\textnormal{ID}}(\bm{x},m)$ is an upper envelope of the subset Lyapunov functions $h(\bm{x},[Nm^{\prime}])$’s. The function $h_{\textnormal{ID}}(\bm{x},m)$ has properties similar to those in Lemma 3, which we state as Lemma 7 and prove in Appendix F.3.
@@
We next introduce the concept of the focus set, which is directly tied to the function $m(\cdot)$ in the Lyapunov function $V$. The focus set is a dynamic subset of arms based on the current system state. Specifically, for any system state $\bm{x}$, the focus set is defined as the set $Nm(\bm{x})]$, where $m(\bm{x})$ is given by
@@
The focus set is introduced because it has several desirable properties that are useful for the analysis. First, under the ID policy, almost all the arms in the focus set, except for $O(\sqrt{N})$ arms, can follow the optimal single-armed policies. Additionally, as the focus set evolves over time, it is almost non-shrinking, and its size is closely related to the value of the function $h_{\textnormal{ID}}(\cdot,\cdot)$. These properties are formalized as Lemmas 8, 9 and 10, which are presented in Appendix G.
@@
where $L_{h}$ is the Lipschitz constant in Lemma 3 and $K_{\text{conf}}$ is the positive constant in Lemma˜8.
@@
The proofs of Lemmas 4 and 5 are provided in Appendix H. It is then straightforward to combine these two lemmas to get Theorem 1.
Appendix F Lemmas and proofs for subset Lyapunov functions
In this section, we prove lemmas for the subset Lyapunov functions $h(\bm{x},D)$ and $h_{\textnormal{ID}}(\bm{x},m)$, defined in (23) and (27). In Section F.1, we prove a preliminary lemma that helps us utilize ˜1. In Section F.2, we prove Lemma 3, which addresses the properties of the subset Lyapunov functions $h(\bm{x},D)$. Finally, in Section F.3, we present and prove Lemma 7, which establishes properties of the function $h_{\textnormal{ID}}(\bm{x},m)$.
F.1 Lemma for utilizing mixing time bound of transition matrices
Recall that in ˜1, we assume that for each arm $i$, the mixing times of the transition matrix $P_{i}$ under the optimal single-armed policy $\bar{\pi}$$^{*}$} is uniformly bounded by a constant $\tau$. In this section, we prove a lemma that allows us to utilize this assumption.
Let $\Xi_{i}$ denote a matrix whose rows are the same vector $\mu_{i}^{*}$, where recall that $\mu_{i}^{*}$ is the stationary distribution of $P_{i}$. Lemma˜6 stated below bounds the infinite series $\sum_{\ell=0}^{\infty}\max_{i\in[N]}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\gamma^{-\ell}$ by a constant independent of $N$.
@@ F.1 Lemma for utilizing mixing time bound of transition matrices
Note that Lemma˜6 is the only place in our proofs where ˜1 is directly invoked; all of our later proofs utilize ˜1 indirectly by invoking Lemma˜6.
Proof of Lemma˜6.
@@
Moreover, by ˜1, it is not hard to see that
@@
Observe that $\tau\geq 1$, $0<1/(2\tau)\leq 1/2$, and $\exp(-a)\leq 1-(1-1/\sqrt{e})a$ for any $0<a\leq 1/2$, so
@@
F.2 Proof of Lemma 3
In this subsection, we prove Lemma˜3, which is about properties of the Lyapunov function $h(\bm{x},D)$:
@@ F.2 Proof of Lemma 3
Lemma˜3 is restated as follows. See 3
In the proof of Lemma˜3, we will frequently use the following equivalent form of $h({\bm{X}},D)$:
@@
By Lemma˜6, $\sum_{\ell=0}^{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}/\gamma^{\ell}$ is finite, so the expression in (36) is also finite. Taking maximum over $g\in\mathcal{G}$, because $\mathcal{G}$ is a finite set, we have
$h(\bm{x},D)=\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\sum_{i\in D}\left\langle(x_{i}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\right\rvert$$\infty$.
@@
Next, we bound the term $\mathbb{E}\left[\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$. Let $\epsilon_{i,t}\in\mathbb{R}^{\mathbb{S}}$ be the random vector given by $\epsilon_{i,t}=X_{i,t+1}-X_{i,t}P_{i}$ for $i\in D$. Then for each state $s\in\mathbb{S}$, $\epsilon_{i,t}(s)$ conditioned on ${\bm{X}}_{t}$ is a Bernoulli distribution with mean 0. Consequently,
[…]
@@
where the inequality in (45) is due to $\sum_{\ell=0}^{\infty}\max_{i\in D}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}/\gamma^{\ell}\leq C_{\tau}$ (Lemma˜6).
@@
F.3 Properties of $h_{\textnormal{ID}}(\cdot,\cdot)$
@@
    where $L_{h}>0$ is the Lipschitz constant given in Lemma˜3.
@@
    where $C_{h}>0$ is the constant given in Lemma˜3.
@@
where (49) is due to the Lipschitz continuity of $h(x,D)$ with respect to $D$, as established in Lemma˜3. We have thus proved (47).
@@
[…]
@@
[…]
@@
Appendix G Lemmas for focus set
In this section, we present and prove three lemmas about properties of the focus set. Recall that for any system state $\bm{x}$, the focus set is defined as the set $Nm(\bm{x})]$, where $m(\bm{x})$ is given by
@@ Appendix G Lemmas for focus set
[…]
@@
[…]
@@
G.1 Proof of Lemma 8 (Majority conformity)
@@
where (G.1) follows from the strict slope of the remaining budget $\overline{C}^{*}([Nm])$ (Lemma˜2). Therefore, choosing $m=m({\bm{X}}_{t})-(\Delta_{t}+M_{c})/(\eta_{c}N)$, we obtain
@@
G.2 Proof of Lemma 9 (Almost non-shrinking)
@@
where $C_{h}>0$ is the positive constant given in Lemma˜3.
@@
We upper bound $\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]$ by coupling ${\bm{X}}_{t+1}$ with a random element ${\bm{X}}_{t+1}^{\prime}$ constructed as follows: Let ${\bm{X}}_{t+1}^{\prime}$ be the system state at step $t+1$ if we were able to set $A_{i,t}=\widehat{A}_{i,t}$ for all $i\in[N]$. From the drift condition of the Lyapunov function $h_{\textnormal{ID}}(\cdot,D)$, as established in Lemma˜7, we obtain:
@@
[…]
@@
where we have applied Lemma˜8 to bound the expression in (66).
@@
where the right-hand side can be further bounded using Lemma˜11, which states that
@@
G.3 Proof of Lemma 10 (Sufficient coverage)
@@
Appendix H Proofs of Lemma 4 and Lemma 5
In this section, we provide two final lemmas, Lemma˜4 and Lemma˜5, which together imply Theorem˜1. We prove these two lemmas in Sections H.1 and H.2, respectively.
H.1 Proof of Lemma 4
@@
where (73) is due to the definition of $N_{t}^{*}$; (74) is due to the bound on $\mathbb{E}[\left(Nm({\bm{X}}_{t})-N_{t}^{*}\right)^{+}]$ in Lemma˜8, and (75) directly follows from the definition of $h_{\textnormal{ID}}(\bm{x},1)$. We further bound the first two expressions in (75) in terms of $V({\bm{X}}_{t})$ as follows:
@@ H.2 Proof of Lemma 5
where the last inequality follows from Lemma˜9. To bound $\mathbb{E}\left[h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))\mid{\bm{X}}_{t}\right]$, observe that by Lemma˜8, all but $O(\sqrt{N})$ arms in $Nm({\bm{X}}_{t})]$ follow the optimal single-armed policies, so the drift condition of $h_{\textnormal{ID}}$ applies to this set of arms. As formalized in Lemma˜11, we can thus show that
@@
[…]
To further bound $h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))$ in (80) in terms of $V({\bm{X}}_{t})$, we apply Lemma˜10 to get:
@@
Averaging over $T$ time steps, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[V({\bm{X}}_{t})\right]\leq\frac{\mathbb{E}\left[V({\bm{X}}_{0})\right]}{\rho_{V}T}+\frac{K_{V}\sqrt{N}}{\rho_{V}}.$ (83)
Taking the limit of $T\to\infty$, we can bound the long-run-averaged expectation of $\mathbb{E}\left[V(\bm{x})\right]$ as
@@
By examining the proof, we can get a more explicit form of the constant $C_{\textnormal{ID}}$ in Theorem 1:
@@
  • If $c_{\max}\geq\alpha_{\min}$, we combine the optimality gap bound in Lemma˜4 with (84) to get
@@
Appendix I Proof of Proposition˜1
Now we briefly outline the proof of Proposition˜1, which states that under the same conditions as Theorem˜1, for any $T\geq 1$, we have
$R^{*}(N,\bm{S}_{0})-\frac{1}{TN}\sum_{t=0}^{T-1}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]\leq\frac{C_{\textnormal{ID}}}{\sqrt{N}}+\frac{C_{\textnormal{finite}}}{T},$ (85)
where $C_{\textnormal{finite}}$ is another positive constant independent of $N$. This follows from two steps:
  • First, it is not hard to adapt the proof of Lemma˜4 in Section˜H.1 to show that
    $R^{*}(N,\bm{S}_{0})-\frac{1}{TN}\sum_{t=0}^{T-1}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$ (86)
    $\qquad\leq\left(1+\frac{2r_{\max}}{L_{h}}\right)\frac{1}{TN}\sum_{t=0}^{T-1}\mathbb{E}[V(\bm{X}_{t})]+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}.$
    Specifically, substituting (77) and (76) into (75) implies that for any $t\geq 0$, we have
    $R^{\textup{rel}}-\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]\leq\left(1+\frac{2r_{\max}}{L_{h}}\right)\frac{1}{N}\mathbb{E}[V(\bm{X}_{t})]+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}.$ (87)
    Averaging the above inequality over $t=0,1,2,\dots T-1$ and applying $R^{*}(N,\bm{S}_{0})\leq R^{\textup{rel}}$, we have (86).
  • Second, we apply (83) in the proof of Lemma˜5 to get
    $R^{*}(N,\bm{S}_{0})-\frac{1}{TN}\sum_{t=0}^{T-1}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$
    $\qquad\leq\left(1+\frac{2r_{\max}}{L_{h}}\right)\left(\frac{\mathbb{E}\left[V({\bm{X}}_{0})\right]}{\rho_{V}TN}+\frac{K_{V}}{\rho_{V}\sqrt{N}}\right)+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}$
    $\qquad=\left(1+\frac{2r_{\max}}{L_{h}}\right)\frac{\mathbb{E}\left[V({\bm{X}}_{0})\right]}{\rho_{V}TN}+\frac{C_{\textnormal{ID}}}{\sqrt{N}}$
    $\qquad\leq\left(1+\frac{2r_{\max}}{L_{h}}\right)\frac{L_{h}}{\rho_{V}T}+\frac{C_{\textnormal{ID}}}{\sqrt{N}}.$ (88)
    where the last inequality is because of $V({\bm{X}}_{0})\leq NL_{h}$ for any system state ${\bm{X}}_{0}$. Letting $C_{\textnormal{finite}}=\left(1+\frac{2r_{\max}}{L_{h}}\right)\frac{L_{h}}{\rho_{V}}$, we get (85).